OpenStudy (anonymous):

250 cm^3 of CO2 weighs 0.44 g while the same volume of a gaseous oxide of Nitrogen weighs 0.46 g at the same temperature and pressure. Find the mass of 1 mole of this oxide of nitrogen. Suggest the molecular formula. Thanks!

OpenStudy (anonymous):

@kausarsalley

OpenStudy (anonymous):

i found the the mass of 1 mole of the oxide nitrogen (Molar mass)... Avogadro's law states that equal volumes of all gases under the same conditions of temperature and pressure contain the same number of moles or molecules.....

OpenStudy (anonymous):

\[n=\frac{ m }{ M }\] n=no. of moles m=mass M=Molar mass(mass of 1 mole)

OpenStudy (anonymous):

using that, \[n(CO_2)=n(nitrogen ~~oxide)\] \[\frac{ 0.44 }{ 12+32 } = \frac{ 0.46 }{ x }\] where x is the Molar mass (mass of 1 mole) of the nitrogen oxide..

OpenStudy (anonymous):

@thomaster do you have an idea??

OpenStudy (anonymous):

@kym02 do you understand what i have done so far??

thomaster (thomaster):

yes @kausarsalley the moles \(\sf CO_2\) = \(\sf\Large\frac{0.44}{44}=\large 0.01~mol\) 0.01 mol of the nitrogen oxide weighs 0.46 gram. So 1 mol of this nitrogen oxide = \(\sf\Large\frac{1}{0.01}\large*0.46\) = 46 g/mol Now what nitrogen oxide has a Mw of 46 g/mol?

OpenStudy (anonymous):

yes...

thomaster (thomaster):

possibilities for nitrogen oxide are: \(\sf \\ \to~NO\\ \to~NO_2\\ \to~N_2O\\ \to~N_2O_3\\ \) There are many more but you probably already passed 46 g/mol

OpenStudy (anonymous):

@kausarsalley yes i do...

OpenStudy (anonymous):

Is the molecular formula NO2?

thomaster (thomaster):

N=14 O=16

OpenStudy (nincompoop):

1 mol of anything is the AMU of that element or the sum of the AMU's of the molecule of compound. to put what thomaster said into proper jargon, an oxide of nitrogen means \[NO_x\] or \[N_xO_x\] the AMU of N is 14.007 and O is 15.999 now let us get teh right combination that satisfies 46g/mol 1 mol N + 1 mol O gives us 30.006 g of NO this means we're missing 16 g more and the best way to achieve that is adding 1 mol of O again. This gives us \[1 mol NO_2 (\frac{ 46.005gNO_2 }{ 1 mol NO_2 })=46.005gNO_2\]

OpenStudy (anonymous):

hmmm thanks !